A 0.60 π‘˜π‘” book slides on a horizontal table. The kinetic friction force on the book has magnitude 1.8 𝑁. (a) How much work is done on the book by friction during a displacement of 3.0 π‘š to the left? (b) The book now slides 3.0 π‘š to the right, returning to its starting point. During this second 3.0 π‘š displacement, how much work is done on the book by friction? (c) What is the total work done on the book by friction during the complete round trip? (d) Based on your answer to part (c), would you say that the friction force is conservative or non-conservative? Explain. 


1
Expert's answer
2021-10-31T18:11:13-0400

(a) The work is


W=fd=5.4 J.W=fd=5.4\text{ J}.

(b) The work is


W=fx=5.4 J.W=fx=5.4\text{ J}.


(c) The total work is


W=f(x+d)=10.8 J.W=f(x+d)=10.8\text{ J}.

(d) Friction is non-conservative because the work depends on the path taken.